OpenStudy (anonymous):

If 8000 dollars is invested in a bank account at an interest rate of 9 per cent per year, find the amount in the bank after 12 years if interest is compounded annually

OpenStudy (wolf1728):

total = principal * (1+rate)^years total = 8,000 * (1.09)^12 total = 8,000 * 2.8126647818 total = 22,501.32 and to check your answer http://www.1728.org/compint.htm